Circular No. 6751

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

     URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1997 T1

S. Nakano, Sumoto, Japan, reports the visual discovery of a comet by Syogo Utsunomiya on Oct. 3. Though Utsunomiya saw no tail, most of the CCD astrometrists below report a tail 4′–12′ long in p.a. ∼ 130^{o}:

1997 UT  α_{2000}  δ_{2000}  m_{1} Observer
Oct. 3.559  23^{h}27^{m}  + 72^{o}42′  10.5 Utsunomiya
5.524  22 40  +71 58  10.5  _{”}
5.65312 22 37 00.^{s}39  +71 53 19′′.3 10.5 Kobayashi
5.65903 22 36 52.53  +71 53 05.9  Kojima
5.66062 22 36 50.40  +71 53 02.3 10.9  _{”}
5.66134 22 36 49.43  +71 53 00.6  Kushida
5.66388 22 36 45.79  +71 52 55.3  _{”}
5.66434 22 36 45.21  +71 52 55.2  Kobayashi
5.66524 22 36 43.95  +71 52 52.1  Kojima
5.66736 22 36 41.08  +71 52 48.0  Kushida
5.67014 22 36 37.19  +71 52 41.6  _{”}
5.67251 22 36 33.98  +71 52 36.1  Kojima
5.67279 22 36 33.63  +71 52 36.6  Kobayashi
5.67292 22 36 33.46  +71 52 35.6  Kushida
5.69529 22 36 02.55  +71 51 46.8 11.2 Ikari
5.69924 22 35 57.22  +71 51 38.4  _{”}
5.70787 22 35 45.34  +71 51 18.3  Kushida
5.71528 22 35 35.27  +71 51 01.9  _{”}
5.72014 22 35 28.63  +71 50 51.0  _{”}
5.76193 22 34 33.59  +71 49 21.2  Galád
5.76336 22 34 31.80  +71 49 18.1  _{”}
5.76421 22 34 30.40  +71 49 16.8  Tichá
5.76461 22 34 29.92  +71 49 15.7 11.2 Galád
5.76525 22 34 28.96  +71 49 14.6  Tichá
5.76608 22 34 27.86  +71 49 12.7 11.0  _{”}

S. Utsunomiya (Azamihara, Minami-Oguni cho, Aso-gun, Kumamoto-ken).

25×150 binoculars. Coma diameter 2′. J. Kobayashi (Kumamoto). Coma diameter 1′ with strong condensation. T. Kojima (Chiyoda). Coma diameter 1′.7 with strong condensation. Y. Kushida (Yatsugatake South Base Observatory). Y. Ikari (Moriyama). J. Tichá and M. Tichý (Kleť). Coma diameter 3′.5. A. Galád and A. Pravda (Modra). Poor reference-star distribution.

1997 October 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green